a plant manager needs a solution of 77% herbicide to be mixed with a solution of 67% herbicide to form 30 liters of 74% solution. How many liters of the 77% solution must he use?
Answer by mangopeeler07(462)

a plant manager needs a solution of 77% herbicide to be mixed with a solution of 67% herbicide to form 30 liters of 74% solution. How many liters of the 77% solution must he use?

x=liters of 77%
30-x=liters of 67%
coefficient=percent

77x+67(30-x)=74(30)------------------------solution of 77% herbicide mixed with a solution of 67% herbicide = 30 liters of 74% solution.


Distribute
77x+2010-67x=2220

Combine like terms
10x+2010=2220

Subtract 2010 from both sides
10x=210

Divide by 10
x=21

x=liters of 77%
30-x=liters of 67%

Answer:
liters of 77%=21
liters of 67%=9